In the laboratory, a general chemistry student measured the pH of a 0.526 M aqueous solution of ethylamine, C2HsNH2 to be 12.193. Use the information she obtained to determine the Kb for this base. Kb(experiment)
Answer #1

pH + pOH = 14

pOH = 14 - 12.193 = 1.807

pOH = - log [OH]

1.807 = - log [OH]

[OH] = 0.015595 M

Ethyl amine is a weak base:

[OH] = (Kb x C)1/2

[OH]2 = (Kb x C)

Kb =  [OH]2/C = 0.015595 x 0.015595 x 0.526 = 1.279 x 10-4